1/2

Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева ,

Ирина Саверьевна Алексеева ,

Ode a Samarcanda Ode für Samarkand Ode to Samarkand
tradukita de Carlo Minnaja tradukita de Manfred Retzlaff tradukita de Malcolm Jones

Argiope-informo: La aspekto de tiu ĉi versio

estas barita. Supozeble la aŭtoro decidis

tion.

Ich singe dir, o Samarkand, Dies Lied. Du bist so jung ge-

blieben. Du bist in aller Welt bekannt. Wer dich gesehn, der muss

dich lieben.

I sing of you, oh, Samarkand’! Your wondrous features win my

praises. And your dear heart - the Registan

- seems young to me and never ages.

Ein Seher hat dich einst ge-

schaut Im Traum vor langer, langer

Zeit, Und Tamerlan hat dich ge-

baut In deiner ganzen Herrlichkeit.

Its steady beating Ulughbek

could sense beneath the constella-

tions. Oh, Samarkand! Immortal pick of mighty Tamerlan’s creations.

Und deine heil’gen Hallen,

Hort Der Kunst, der Wissenschaft

und Lehre, Jahrhunderte schon stehn sie

dort. Als ob kaum Zeit vergangen

wäre.

Madrasses are your sacred schools, which cultivate the arts and lear-

ning. Their beauty overcomes time’s ru-

les by centuries into moments turning.

Hier hat der Dichter Navoí Von Liebe und von Kraft ge-

sungen, Und seine Lieder sind noch

nie In diesem Lande ganz verk-

lungen.

Here the great Navoi’s verse, so full of love and strong emotion, is cherished by your populace, made happy by its inspiration.

Der Minarette Kuppeln hier, Sie leuchten tags wie nachts

die Sterne. Mit ihrer Ornamente Zier Sie grüßen jeden schon von

ferne.

The domes of minarets stand

proud - like day-time stars in clear sky shi-

ning: so pure the blue, without a cloud and fine the spires in art combi-

ning.

…  …  …

2/2

O Samarkand, der Weisheit

Stadt! Hier stehen Kirchen und Mo-

scheen, Und wer dich liebt, der sagt,

er hat Nichts Schön’res auf der Welt

gesehen.

Oh, Samarkand! You’re wise

through age, with mosques and churches closely

neighboured. For those who love you, the world’s

stage has no town better nor more fa-

voured.

O Samarkand, blüh’ und ge-

deih’, Dein Ruhm sich überall ver-

breite, Und ewig unvergänglich sei, Auch in der Zukunft uns be-

gleite!

Oh, Samarkand! Live long and

thrive, be glorified by each good action! And while eternity’s your prize, be here our lasting benefaction.

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en itala de Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов .

– Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_Minnaja .

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Germanan de Manfred Retzlaff (∗1938-11-04).

Arg-860-1735 (2013-04-03 23:31:01)

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Anglan de Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEA-jarlibro de la jaro 2014 sur paĝo 126.

http://www.poezio.net/version?poem-id=860&version-id=,2206,1735,2037